OpenStudy (amistre64):

Let m,n,p be any integers. If [(m+n) and (n+p) are even integers], then [(m+p) is even].

OpenStudy (amistre64):

another proof :)

OpenStudy (amistre64):

What I wrote in 10 pages or less was: Spose m,n,p are even integers; 2m + 2n = 2(m+n), satisfies the definition of an even integer. Since the addition of any arbitrary even integers is even, then all the results are true. Spose m,n,p are odd integers; 2m+1 + 2n+1 = 2(m+n+1), satisfies the definition of an even integer. Since the addition of any arbitrary odd integers is even, then all the results are true. Spose m,n are odd integers, and p is any arbitrary integer; 2m + 2n+1 = 2(m+n)+1, makes the premise false; therefore the truth value is undetermined.

OpenStudy (amistre64):

which i am sure is unsound

OpenStudy (amistre64):

I think id only need the "if even then even" statement instead of trying to prove the round in a shape

OpenStudy (amistre64):

...to prove that round is a shape . youknow; these would be funnier if i could spell them correctly

OpenStudy (anonymous):

\[m+n=2k+1,n+p=2j+1\] \[m+n+n+p=2k+1+2j+1=2k+2j+2\] \[m+2n+p=2k+2j+2\] \[m+p=2j+2k+2-3n=2(k+j+1-n)\] is even

OpenStudy (anonymous):

man did i mess that up. it says "even" and i did it as if odd! but even is the same

OpenStudy (amistre64):

:) now that odd is proven ... lol

OpenStudy (anonymous):

\[m+n=2k,n+p=2j\] \[m+n+n+p=2k+2j\] \[m+p=2j+2k-2n\] there

OpenStudy (anonymous):

if i don't get to work...

OpenStudy (jamesj):

So, argument by contradiction again. Suppose that m + n and n + p are even AND m + p is odd. Then either m or p is odd, but not both. Suppose first that m is odd and p is even. If m + n is even, then n is is odd. But if n is odd and p is even, then p + n is odd. But p + n is even. Hence it must be that m is even and p is odd. But if p is odd and p + n is even, then n is odd. But n is odd and m is even, then m + n is odd. But m + n is even. Contradiction. Thus if m + n and n + p are even, then m + p is even.

OpenStudy (amistre64):

Yeah, no matter how i look at it is just seems to be too wordy. But then a proof is only as good as those who agree on it :)

OpenStudy (jamesj):

No. It's good if it's correct. And it is. :-) Read it!

OpenStudy (amistre64):

i read it, and i agree with it, and i like it

OpenStudy (amistre64):

m+n = 2k n+p = 2j ----------------- m+n+n+p = 2k + 2j m+p+2n = 2(k+j) m+p = 2(k+j) - 2n m+p = 2(k+j-n)

OpenStudy (amistre64):

oh .... thats satellites proof

OpenStudy (jamesj):

That's good too.